OpenStudy (anonymous):

Need help to factorise 2m^12-1/8?

OpenStudy (unklerhaukus):

\[2m^{12}-\tfrac18\] its a difference of cubes

OpenStudy (anonymous):

how is this n cube?

OpenStudy (unklerhaukus):

do you know the cube root of eight?

OpenStudy (anonymous):

i think its 2?

OpenStudy (unklerhaukus):

yes so the cube root of 1/8 is ?

OpenStudy (anonymous):

1/2?

OpenStudy (unklerhaukus):

that's right

OpenStudy (unklerhaukus):

now can you take the cube root of \( m^{12}\) ? remember that \[\sqrt[3]x=x^{1/3}\]and \[(y^a)^b=y^{ab}\]

OpenStudy (anonymous):

12root of m, i dont know how to make the root sign. sorry

OpenStudy (unklerhaukus):

[alt][v] √

OpenStudy (unklerhaukus):

for cube root ^3√

OpenStudy (unklerhaukus):

so what is ^3√(m^12) = (m^12)^(1/3)

OpenStudy (anonymous):

give me a moment to think about this, its long time since i did any cube functions

OpenStudy (anonymous):

A bit simpler? \[2m^{12}-\frac{ 1 }{ 8 }=2(m^{12}-\frac{ 1 }{ 16 })=2(m^6+\frac{ 1 }{ 4 })(m^6-\frac{ 1 }{ 4 })=\]\[=2(m^6+\frac{ 1 }{ 4 })(m^3+\frac{ 1 }{ 2 })(m^3-\frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

2m^12-1/8 = 2 ( m^12 - 1/16) then you get: 2 ( m^12 - 1/16) = 2 (m^6 1/4) (m^6 - 1/4) = 2 (m^6 1/4) (m^3 1/2) (m^3 - 1/2) this is what i originally got but was unsure because i did this stuff along time ago, so i just wanted to check with you guys cause my confidence in maths is not there yet

OpenStudy (anonymous):

That is correct

OpenStudy (anonymous):

CarlosGP : how do you get it to type your factors like that, i have no clue

OpenStudy (anonymous):

use the button:\[\sum_{}^{}Equation\] (bottom left of the text editor)

OpenStudy (unklerhaukus):

``` \[\frac{a+b}{d}\] ```

OpenStudy (anonymous):

thanks you guys

OpenStudy (anonymous):

u r welcome